# PhysVid: Physics Aware Local Conditioning for Generative Video Models

### 27 Mar 2026 v1

@@ 4.1 Setup
We use the WISA [48] data set that contains a diverse collection of more than $80$ thousand videos related to various physical phenomena observed in the world. Following the configuration in Wan [47], we remove videos less than $5$ seconds long and divide the remaining videos into $5$ second clips sampled at $832\times 480$ at $16$ frames per second, leading to a total of $81$ frames per video. This results in approximately $53$ thousand video samples. Note that while WISA provides detailed physics annotations for each of their training samples, we do not utilize them in our approach, relying instead on learning this information purely from the training videos. This is because these annotations are at a global level and do not focus on physical phenomena observed within smaller temporal segments of the video. Furthermore, due to the division of long videos into $5$ second clips, these global annotations become misaligned with the video clips and may lead to noisy conditioning if used directly. This strategy also has the added benefit of making our approach generalizable to other datasets that do not contain such explicit information about physical phenomena. Subsequently, we generate our own chunk-level physics-based prompts using the method described in Section 3 as well as a global text caption for the entire video. We use VideoLLama3-7B [64] for this task. First, we generate a global caption for the entire $5$ second video clip. Thereafter, the video is divided into $7$ temporally contiguous chunks of frames of approximately $0.7$ seconds each. Next, we generate a segment-level prompt using the video chunk along with the global caption as input to the VLM, processing each chunk separately in this manner. Examples of the generated prompts are provided in Appendix 8.
@@
We introduce chunk-aware cross-attention layers in each transformer block in a pretrained Wan2.1 [47] model with $1.3$ billion-base parameters. The entire architecture is trained in two stages. First, the base model is frozen, and only the newly added modules are trained for $1000$ steps. Once they have stabilized, the base layers are unfrozen, and the entire model is further trained for additional $2000$ steps. We use $4$ GPUs for this task, leading to an effective batch size of $64$ samples per step. To generate annotations during inference as discussed in Sec. 3, we utilize VideoLLama3-7B as a language model.
@@ 4.2 Evaluation benchmarks
We use two recently proposed and widely adopted benchmarks, VideoPhy [3] and VideoPhy2 [4], to evaluate the physical accuracy of the generated videos. VideoPhy consists of $344$ manually curated captions in three different categories of physical interactions, namely, “solid-solid”, “solid-fluid”, and “fluid-fluid”. Similarly, VideoPhy2 uses a larger test set with $590$ captions providing coverage over a diverse set of real-world physical phenomena. Their data is divided into two main action categories: “object interactions” and “sports and physical activities”. Both these benchmarks include an additional category comprising of manually labeled “hard” examples. In addition, they provide an automatic evaluation model that is trained on human annotations of the generated videos. As a result, their evaluation scores are correlated with human judgment of physical correctness. We follow the same scoring mechanism as outlined in the respective original works. All results are reported as the mean of the benchmark scores for $5$ different evaluation sets generated with different random seeds. We present the results next.
@@ 4.3.1 Quantitative results
As shown in Tab. 1, PhysVid performs, respectively, $\approx 5\%$ and $\approx 8\%$ better than the corresponding baselines on the physical commonsense score. As Fig. 6 shows, this improvement is consistent across both “object interaction” and “sports and physical activities” subcategories of the benchmark, as well as on the captions that are categorized by the benchmark as hard to generate accurately. Compared to VideoPhy, the improvements of our method are significantly less pronounced on VideoPhy2. The reason behind this could be the difference in the score calculation method between the two evaluation approaches. Specifically, VideoPhy2 uses a categorical rating system in contrast to VideoPhy, which uses a rating on a continuous scale of 0 and $1$, followed by hard thresholding [3, 48].

### 30 Mar 2026 v2

@@ 4.1 Setup
We use the WISA [48] data set that contains a diverse collection of more than $80$ thousand videos related to various physical phenomena observed in the world. Following the configuration in Wan [47], we remove videos less than 55 seconds long and divide the remaining videos into 55 second clips sampled at $832\times 480$ at $16$ frames per second, leading to a total of $81$ frames per video. This results in approximately $53$ thousand video samples. Note that while WISA provides detailed physics annotations for each of their training samples, we do not utilize them in our approach, relying instead on learning this information purely from the training videos. This is because these annotations are at a global level and do not focus on physical phenomena observed within smaller temporal segments of the video. Furthermore, due to the division of long videos into 55 second clips, these global annotations become misaligned with the video clips and may lead to noisy conditioning if used directly. This strategy also has the added benefit of making our approach generalizable to other datasets that do not contain such explicit information about physical phenomena. Subsequently, we generate our own chunk-level physics-based prompts using the method described in Section 3 as well as a global text caption for the entire video. We use VideoLLama3-7B [64] for this task. First, we generate a global caption for the entire 55 second video clip. Thereafter, the video is divided into 77 temporally contiguous chunks of frames of approximately $0.7$ seconds each. Next, we generate a segment-level prompt using the video chunk along with the global caption as input to the VLM, processing each chunk separately in this manner. Examples of the generated prompts are provided in Appendix 8.
@@
We introduce chunk-aware cross-attention layers in each transformer block in a pretrained Wan2.1 [47] model with $1.3$ billion-base parameters. The entire architecture is trained in two stages. First, the base model is frozen, and only the newly added modules are trained for $1000$ steps. Once they have stabilized, the base layers are unfrozen, and the entire model is further trained for additional $2000$ steps. We use 44 GPUs for this task, leading to an effective batch size of $64$ samples per step. To generate annotations during inference as discussed in Sec. 3, we utilize VideoLLama3-7B as a language model.
@@ 4.2 Evaluation benchmarks
We use two recently proposed and widely adopted benchmarks, VideoPhy [3] and VideoPhy2 [4], to evaluate the physical accuracy of the generated videos. VideoPhy consists of $344$ manually curated captions in three different categories of physical interactions, namely, “solid-solid”, “solid-fluid”, and “fluid-fluid”. Similarly, VideoPhy2 uses a larger test set with $590$ captions providing coverage over a diverse set of real-world physical phenomena. Their data is divided into two main action categories: “object interactions” and “sports and physical activities”. Both these benchmarks include an additional category comprising of manually labeled “hard” examples. In addition, they provide an automatic evaluation model that is trained on human annotations of the generated videos. As a result, their evaluation scores are correlated with human judgment of physical correctness. We follow the same scoring mechanism as outlined in the respective original works. All results are reported as the mean of the benchmark scores for 55 different evaluation sets generated with different random seeds. We present the results next.
@@ 4.3.1 Quantitative results
As shown in Tab. 1, PhysVid performs, respectively, $\approx 5\%$ and $\approx 8\%$ better than the corresponding baselines on the physical commonsense score. As Fig. 6 shows, this improvement is consistent across both “object interaction” and “sports and physical activities” subcategories of the benchmark, as well as on the captions that are categorized by the benchmark as hard to generate accurately. Compared to VideoPhy, the improvements of our method are significantly less pronounced on VideoPhy2. The reason behind this could be the difference in the score calculation method between the two evaluation approaches. Specifically, VideoPhy2 uses a categorical rating system in contrast to VideoPhy, which uses a rating on a continuous scale of 0 and 11, followed by hard thresholding [3, 48].